# Evaluation Should Not Ignore Variation: On the Impact of Reference Set Choice on Summarization Metrics

### v1

@@ 4.3 System Ranking Depends on the Reference(s) for n-gram-based Metrics
While we observed variability at the instance level, summarization metrics are typically designed to evaluate models across entire datasets, rather than individual instances. In this section, we investigate to what degree standard summarization metrics can handle the variability observed in human-written references when ranking summarization systems.
@@
We sample $k$ humans-written references ($k\in[1,N]$, where $N$ is the number of references for each document) from all available references for each instance. We then score the outputs of each summarization system using the same set of references. Given $M$ systems $S_{a}$, $S_{b}$, …, $S_{M}$, the metric induces a ranking $S_{a}\succ S_{b}\succ\dots\succ S_{M}$. This process is repeated $100$ times, yielding $100$ resulting rankings. We compute the pairwise Kendall rank correlation coefficient Kendall (1938) between such ranks. High correlation indicates that different sets of references lead to similar model ordering, even when different sets of references are used. Figure 5 reports the average correlation for pairs of ranks for each dataset and metric, from using a $k$ human-written summaries as references. ROUGEmax is shown in Figure 5, and ROUGEavg is reported in Figure 11 in Appendix F.
@@
Evaluating with a single reference is common in summarization, as most datasets provide only one human-written summary. Figure 5 (looking at $k=1$) shows the stability of different metrics with a single reference across the three datasets. We find that BLEU and ROUGE have very weak to moderate correlation between ranks across different references. In other words, using two different sets of plausible references would likely lead to different conclusions on relative model performance. We also notice a large variability among the individual pairs of rankings, with some showing negative correlation (refer to Table 4 in Appendix F for results on individual metrics and datasets).
In contrast, more semantically-oriented metrics show a greater stability. For SummEval, BLEURT shows the highest correlation between ranks, followed by METEOR and BERTScore. BLEURT and METEOR confirm their stability on GUMSum when ranking the LLM outputs. Other metrics (including BERTScore) show low or no correlation on GUMSum, with the exception of ROUGE-1. In all cases, metrics show much higher stability on DUC, for which all average correlations are above $0.7$. We speculate that the high stability might be due to an artifact introduced by the short length of the summaries, required by the guidelines.
@@
When scoring model outputs against a set of $k>1$ randomly sampled references, we observe that the correlation between rankings obtained with different human-written references generally improves when increasing the number of references. This increased stability is expected and in line with similar findings that associate a larger number of references with a higher correlation with humans Lin (2004).
However, the stability varies by metric. ROUGE (especially ROUGEmax) and BLEU tend to have lower correlation between ranks than other metrics. As an example, the ROUGEmax scores require $5$-$10$ references to reach a level of stability that is comparable to that of BERTScore on SummEval with a single reference. ROUGEavg has a better stability than ROUGEmax, especially with a larger set of references. For example, on SummEval, ROUGE-Lavg has higher stability than BERTScore for $k>3$, while on GUMSum, ROUGE-2avg is the second most stable metric for $k>3$. On all datasets, BLEURT and METEOR remain stable even with a single reference, with METEOR showing remarkable stability despite its simplicity.
In general, trends on SummEval are clearer and simpler to interpret than the other two datasets. We speculate that this is due to the larger number of models used ($16$ pre-LLM models+$4$ LLMs on SummEval vs $4$ LLMs on GUMSum and DUC). BLEURT, METEOR, and BERTScore show the highest stability, while n-gram-based metrics show low to average correlation between ranks even when multiple references are used. The cases of GUMSum and DUC2004 are more complex to interpret and might be less meaningful given fewer model outputs (i.e., only four LLM outputs). For GUMSum, BLEURT continues to show high inter-rank correlation, with METEOR being the second most stable. BERTScore, on the other hand, shows poor stability. Similar to the case with $k=1$, on DUC2004, all metrics show high stability, likely due to summaries being very short, as dictated by the guidelines.
@@ 4.4 Correlation with Human Judgments
In addition to stability, automatic metrics should correlate with human judgments. We compute correlations for SummEval and GUMSum, for which we have human judgments, [^5] For SummEval, we use the $16$ models studied by Fabbri et al. (2021); for GUMSum, the four LLMs. at the instance and system level as the number of references $k$ increases. [^6] For example, when considering two references, we consider the sets ($\binom{N}{2}$) of human-written references, where $N$ is the total number of references. We compute the scores using such references as gold standard, and report the mean.
@@
Figure 6 reports the instance-level correlation for SummEval (top) and GUMSum (bottom), respectively, versus number of references. We show ROUGEmax; corresponding figures using ROUGEavg are in Appendix G.
@@
System-level correlation is generally higher than instance-level correlation on SummEval; however, many criteria still show weak to moderate correlation when one or very few references are included. In most cases, such correlation tends to improve with the number of references. This is not the case for ROUGEmax, especially when considering consistency. The full results are provided in Figure 13 in the Appendix G. GUMSum is excluded from this analysis due to the small number of systems available.
@@ 5 Conclusions
In this work, we investigate how reference sets impacts the reliability of reference-based summarization metrics. Our analysis across three multi-reference datasets—SummEval, GUMSum, and DUC2004—reveals that token-matching metrics such as ROUGE and BLEU are highly sensitive to the reference(s). This sensitivity leads to instability in system rankings, particularly when only a small number of references are available. We thus recommend avoiding such metrics, complementing previous work Schmidtova et al. (2024).
We find that increasing the number of reference summaries consistently improves both the stability of metric scores and their alignment with human judgments. In these conditions, n-gram-based metrics such as ROUGE-N become more reliable. Moreover, ROUGEavg should typically be preferred over ROUGEmax. However, when few references are available—especially in domains with high output variability—model-based metrics should be preferred, as they offer greater robustness despite their higher computational costs.
Our findings highlight the need to incorporate reference set variation into evaluation frameworks and suggest that future metric development should explicitly account for this dimension. We thus advocate for the creation of larger, more diverse multi-reference datasets, as well as for metric designs that are inherently robust to reference variability. Such efforts will be key to ensuring fairer and more reliable and human-aligned evaluation practices in summarization in the LLM era and beyond.
@@
Firstly, although we focus on multi-reference datasets—SummEval, DUC2004, and GUMSum—such datasets remain relatively small. This reflects a broader limitation in current meta-evaluation practices, where multi-reference resources are the norm despite their limited scale.
Secondly, our analysis primarily targets standard evaluation criteria such as coherence, consistency, fluency, and relevance. While these are widely adopted and established, they do not capture all the nuances of summary quality, especially when the source texts are genre-diverse. More fine-grained human annotations and task-specific dimensions (e.g., factuality for news, stance for opinion pieces) would allow a deeper understanding of metric behavior under reference variation.
Thirdly, while we assess metric robustness using four systems (all of which are LLMs as opposed to the case for SummEval, where $16$ pre-LLM supervised systems are available) on the GUMSum and DUC2004 datasets, the relatively small number of models limits the generalization of the conclusions we can draw about system-level stability. Future work should include a larger and more diverse set of systems to better assess generalizability.
Lastly, despite our efforts to use contemporary LLMs, there remains a potential risk of data contamination. Since some of these datasets may have been seen during pretraining, this could affect both the outputs generated by LLMs and their evaluation scores. While we do not observe signs of memorization, we acknowledge that further controlled experiments are necessary to rigorously assess this risk.
@@
Acknowledgments
We would like to thank the members of the MaiNLP lab for their valuable feedback, especially to Andreas Säuberli, Diego Frassinelli, Shijia Zhou, Soh-Eun (Ryan) Shim, and Domenico De Cristofaro. We also thank Pingjun Hong, Yuchen Mao, and Sebastian Loftus for their annotations. We recognize the support for Yang Janet Liu and Barbara Plank through the ERC Consolidator Grant DIALECT 101043235.
@@ Appendix A LLM-generated Summaries
For each source, we generate a single summary using each of the four LLMs. For SummEval, the sources corresponding to the $100$ multi-reference summaries are taken into account. For DUC2004 Task 1, we generate summaries for the whole dataset ($489$ instances). For GUMSum, we focus on the dev and test set (in total $48$ instances). Thus, we generate a total of $2548$ summaries.
@@ Appendix B Reference-based Metrics
We use the original Google version of BLEURT. [^13] https://github.com/google-research/bleurt We used the recommended checkpoint,141414available at https://storage.googleapis.com/bleurt-oss-21/BLEURT-20.zip which we did not fine-tune.
Appendix C Variability in Humans and LLMs
@@
Appendix D Quality of Scraped References
Figure 8: Models and CNN/DM reference compared to $10$ crowd-sourced references in SummEval. SummEval models are in lighter blue, LLMs in darker blue. The red line indicates the CNN/DM reference score, and the black dotted line represents a random baseline using a different news article as the hypothesis.
@@ Appendix D Quality of Scraped References
Figure 8 reports the scores of the original CNN/DM references (red line) and those of the outputs of the summarization systems (blue bars). We also report a random baseline (dotted line) in which a summary of a different document randomly sampled from the collection is used as hypothesis.
Notably, the original CNN/DM references perform worse than all model outputs in all cases but one when evaluated using n-gram-matching metrics. When using BERTScore, the CNN/DM receives a score close to that of the outputs from LLMs. BLEURT rates the original reference higher than SummEval system outputs (except for GPT-2), but still lower than all LLM-generated outputs. These observations corroborate previous concerns on reference quality for summarization, especially when used to score high-quality outputs Goyal et al. (2023).
Appendix E Instance-level variation
Figures 9 and 10 contain the histogram of the instance-level variability for GUMSum and DUC respectively.
@@ Appendix E Instance-level variation
Appendix F Model Ranking
@@
Table 4: Rank stability with a single reference. We ranked systems $100$ times and compute the Kendall tau correlation among such rankings. R, MTR, BS, and BLRT are short for ROUGE, METEOR, BERTScore, and BLEURT respectively.
@@
Appendix G System-level Correlation
Figure 12 shows the instance-level correlation for SummEval (top) and GUMSum (bottom) using ROUGEavg.
Figure 13 show the system-level correlation on SummEval with ROUGEmax (top) and ROUGEavg (bottom).

### v2

@@ 4.3 System Ranking Depends on the Reference(s) for N-gram-based Metrics
While we observed variability at the instance level, summarization metrics are typically designed to evaluate models across datasets, rather than individual instances. In this section, we investigate to what degree standard summarization metrics can handle the variability observed in human-written references when ranking summarization systems.
@@
We sample $k$ humans-written references ($k\in[1,N]$, where $N$ is the number of references for each document) from all available references for each instance. We then score the outputs of each summarization system using the same set of references. Given $M$ systems $S_{a}$, $S_{b}$, …, $S_{M}$, the metric induces a ranking $S_{a}\succ S_{b}\succ\dots\succ S_{M}$. This process is repeated 100100100 times, yielding 100100100 rankings. We compute the pairwise Kendall rank correlation coefficient Kendall (1938) between such ranks. High correlation indicates that models are similarly ordered, even when different sets of references are used. Figure 6 reports the average correlation for pairs of ranks for each dataset and metric, from using $k$ human-written summaries as references. ROUGEmax is shown in Figure 6, and ROUGEavg is reported in Figure 12 in Appendix G.
@@
Evaluating with a single reference is common in summarization, as most datasets provide only one human-written summary. Figure 6 (looking at $k=1$) shows the stability of different metrics with a single reference across the three datasets. We find that BLEU and ROUGE have very weak to moderate correlation between ranks across different references. In other words, using two different sets of plausible references would likely lead to different conclusions on relative model performance. We also notice a large variability among the individual pairs of rankings, with some showing negative correlation (refer to Table 4 in Appendix G for results on individual metrics and datasets).
In contrast, more semantically-oriented metrics show greater stability. For SummEval, BLEURT shows the highest correlation between ranks, followed by METEOR and BERTScore. BLEURT and METEOR confirm their stability on GUMSum when ranking the LLM outputs. Other metrics (including BERTScore) show low or no correlation on GUMSum, with the exception of ROUGE-1. In all cases, metrics show much higher stability on DUC, for which all average correlations are above 0.70.70.7. We speculate that high stability might stem from an artifact introduced by the short summary length required by the guidelines.
@@
When scoring model outputs against a set of $k>1$ randomly sampled references, we observe that the correlation between rankings obtained with different human-written references generally improves with an increased number of references. This increased stability is expected and in line with similar findings that associate a larger number of references with a higher correlation with humans Lin (2004).
However, the stability varies by metric. ROUGE (especially ROUGEmax) and BLEU tend to have low correlation between ranks. As an example, the ROUGEmax scores require 555-101010 references to reach a level of stability that is comparable to that of BERTScore on SummEval with a single reference. ROUGEavg has a better stability than ROUGEmax, especially with a larger set of references. For example, on SummEval, ROUGE-Lavg has higher stability than BERTScore for $k>3$, while on GUMSum, ROUGE-2avg is the second most stable metric for $k>3$. On all datasets, BLEURT and METEOR remain stable even with a single reference, with METEOR showing stability despite its simplicity.
In general, trends on SummEval are clearer and simpler to interpret than the other two datasets. We speculate that this is due to the larger number of models used (161616 pre-LLM models+444 LLMs on SummEval vs 444 LLMs on GUMSum and DUC). BLEURT, METEOR, and BERTScore show the highest stability, while n-gram-based metrics show low to average correlation between ranks even when multiple references are used. The cases of GUMSum and DUC2004 are more complex to interpret and might be less meaningful given fewer model outputs (i.e., only four LLM outputs). For GUMSum, BLEURT continues to show high inter-rank correlation, with METEOR being the second most stable. BERTScore, on the other hand, shows poor stability. Similar to the case with $k=1$, on DUC2004, all metrics show high stability, likely due to summaries being very short, as dictated by the guidelines.
@@ 4.4 Correlation with Human Judgments
In addition to stability, automatic metrics should correlate with humans. We compute correlations for SummEval and GUMSum, for which we have human judgments, [^4] For SummEval, we use the 161616 models studied by Fabbri et al. (2021); for GUMSum, the four LLMs. at the instance and system level as the number of references $k$ increases. [^5] For example, when considering two references, we consider the sets ($\binom{N}{2}$) of human-written references, where $N$ is the total number of references. We compute the scores using such references as gold standard, and report the mean.
@@
Figure 7 reports the instance-level correlation for SummEval (top) and GUMSum (bottom), respectively, versus number of references. We show ROUGEmax; corresponding figures using ROUGEavg are in Appendix H.
@@
System-level correlation is generally higher than instance-level correlation on SummEval; however, many criteria still show weak to moderate correlation when one or very few references are included. In most cases, such correlation tends to improve with the number of references. This is not the case for ROUGEmax, especially when considering consistency. The full results are provided in Figure 14 in the Appendix H. GUMSum is excluded from this analysis due to the small number of systems available.
@@ 5 Conclusions
In this work, we investigate how reference sets impact the reliability of reference-based summarization metrics. Our analysis across three multi-reference datasets reveals that, despite their popularity, token-matching metrics such as ROUGE are highly sensitive to the reference(s). This sensitivity leads to instability in system rankings, particularly when only a small number of references are available, which is typical in summarization datasets. In these situations, we thus recommend avoiding such metrics, echoing earlier calls for caution Schmidtova et al. (2024), in favor of model-based or reference-free alternatives.
We find that increasing the number of reference summaries consistently improves both the stability of metric scores and their alignment with human judgments. In these conditions, n-gram-based metrics such as ROUGE-N become more reliable.
Our findings highlight the need to incorporate reference set variation into evaluation frameworks. Future metric development should explicitly account for this variation. We thus advocate for the creation of larger, more diverse multi-reference datasets, as well as for metric designs that are inherently robust to reference variability. Such efforts will be key to ensuring fairer and more reliable and human-aligned evaluation practices in summarization in the LLM era and beyond.
@@
Although we focus on multi-reference datasets—SummEval, DUC2004, and GUMSum—such datasets remain relatively small. This reflects a broader limitation in current meta-evaluation practices, where multi-reference resources are the norm despite their limited scale.
Our analysis primarily targets standard evaluation criteria such as coherence, consistency, fluency, and relevance. While these are widely adopted and established, they do not capture all the nuances of summary quality, especially when the source texts are genre-diverse. More fine-grained human annotations and task-specific dimensions (e.g., factuality for news, stance for opinion pieces) would allow a deeper understanding of metric behavior under reference variation.
While we assess metric robustness using four systems (all of which are LLMs as opposed to the case for SummEval, where 161616 pre-LLM supervised systems are available) on the GUMSum and DUC2004 datasets, the relatively small number of models limits the generalization of the conclusions we can draw about system-level stability. Future work should include a larger and more diverse set of systems to better assess generalizability.
Moreover, using reference-based metrics for LLMs outputs (and generally, very high-quality outputs) has been questioned, especially when considering low-quality (e.g., scraped) references. While we focus on high-quality human-written references especially collected and checked for quality, we find that LLM-outputs are scored higher than references in our human evaluation campaign. We want to point out that the paper does not advocate to use reference-based metrics in such context; rather, it aims at shading lights on the limitation of the current evaluation practices, including the use of references in the cases in which the outputs might have higher quality, and its impact on stability and correlation. We also advocate for more research on the similarities and differences between LLM- and human-written summaries, to understand to which extent the use of LLM output as references could improve evaluation or rather introduce unwanted and largely unknown biases.
Lastly, given our efforts to use contemporary LLMs, there remains a potential risk of data contamination. Since some of these datasets may have been seen during pretraining, this could affect both the outputs generated by LLMs and their evaluation scores. While we do not observe signs of memorization, we acknowledge that further controlled experiments are necessary to rigorously assess this risk.
@@ Appendix A LLM-generated Summaries
For each source, we generate a single summary using each of the four LLMs. For SummEval, the sources corresponding to the 100100100 multi-reference summaries are taken into account. For DUC2004 Task 1, we generate summaries for the whole dataset (489489489 instances). For GUMSum, we focus on the dev and test set (in total 484848 instances). Thus, we generate a total of 254825482548 summaries. We comply with the license of the existing datasets. For newly collected model outputs and human judgments, we follow the license of the corresponding underlying datasets. The codebase will be made publicly available upon publication.
@@ Appendix B Reference-based Metrics
We use the original Google version of BLEURT. [^12] https://github.com/google-research/bleurt We used the recommended checkpoint,131313available at https://storage.googleapis.com/bleurt-oss-21/BLEURT-20.zip which we did not fine-tune.
Appendix C Use of Reference-based Metrics
We survey papers published at ACL, EMNLP, and INLG from 2023 and 2024. Given the list of long and short accepted papers, we collected papers matching the keyword summar* in their title. After filtering out papers that are not about summarization research (e.g., summarizing the state of the art for a different topic), for each paper we checked whether one of our chosen metrics had been used for evaluation. Finally, we point out that some of the surveyed papers do not perform experimental work and thus do not evaluate model outputs (e.g., paper studying human evaluation); the reported percentages are thus slightly underestimated.
Appendix D Variability in Humans and LLMs
@@
Appendix E Quality of Scraped References
Figure 9: Models and CNN/DM reference compared to 101010 crowd-sourced references in SummEval. SummEval models are in lighter blue, LLMs in darker blue. The red line indicates the CNN/DM reference score, and the black dotted line represents a random baseline using a different news article as the hypothesis.
@@ Appendix E Quality of Scraped References
Figure 9 reports the scores of the original CNN/DM references (red line) and those of the outputs of the systems (blue bars). We also report a random baseline (dotted line) in which a summary of a different document randomly sampled from the collection is used as hypothesis.
Notably, the original CNN/DM references perform worse than all model outputs in all cases but one when evaluated using n-gram-matching metrics. When using BERTScore, CNN/DM receives a score close to that of the outputs from LLMs. BLEURT rates the original reference higher than SummEval system outputs (except for GPT-2), but still lower than all LLM-generated outputs. These observations corroborate previous concerns on reference summary quality, especially when used to score high-quality outputs Goyal et al. (2023).
Appendix F Instance-level variation
Figure 10 and Figure 11 contain the histogram of the instance-level variability for GUMSum and DUC respectively.
@@ Appendix F Instance-level variation
Appendix G Model Ranking
@@
Table 4: Rank stability with a single reference. We ranked systems 100100100 times and compute the Kendall tau correlation among such rankings. R, MTR, BS, and BLRT are short for ROUGE, METEOR, BERTScore, and BLEURT respectively.
@@
Appendix H System-level Correlation
Figure 13 shows the instance-level correlation for SummEval (top) and GUMSum (bottom) using ROUGEavg. Figure 14 show the system-level correlation on SummEval with ROUGEmax (top) and ROUGEavg (bottom).